OpenStudy (mtwizzle420):

For questions 9-10 find the value of x to the nearest tenth using your calculator.

OpenStudy (mtwizzle420):



OpenStudy (mtwizzle420):

@Mr.Man420 can you help me? :)

OpenStudy (anonymous):

tan42=x/5 x=5*tan42 x=4.5

OpenStudy (mtwizzle420):

Thank you! I have one more can you help me ?:)